OpenStudy (anonymous):

using loarithms solve:the population of a city is growing at a rate proportional to its population. the population 20 years ago was 100 000 and today it is 150 000. find the population 20 years from now

OpenStudy (anonymous):

100 000(r)^20 = 150 000 r^20 = 1.5 r = \[\sqrt[20]{1.5}\]

OpenStudy (anonymous):

hi there

OpenStudy (anonymous):

how are you?

OpenStudy (anonymous):

not good stuck with my test.

OpenStudy (anonymous):

I was just wondering if I have to use this formula P=Po e^+-kt